# PIKE-RAG: sPecIalized KnowledgE and Rationale Augmented Generation

### v2

@@ Abstract
Despite notable advancements in Retrieval-Augmented Generation (RAG) systems that expand large language model (LLM) capabilities through external retrieval, these systems often struggle to meet the complex and diverse needs of real-world industrial applications. The reliance on retrieval alone proves insufficient for extracting deep, domain-specific knowledge performing in logical reasoning from specialized corpora. To address this, we introduce sPecIalized KnowledgE and Rationale Augmentation Generation (PIKE-RAG), focusing on extracting, understanding, and applying specialized knowledge, while constructing coherent rationale to incrementally steer LLMs toward accurate responses. Recognizing the diverse challenges of industrial tasks, we introduce a new paradigm that classifies tasks based on their complexity in knowledge extraction and application, allowing for a systematic evaluation of RAG systems’ problem-solving capabilities. This strategic approach offers a roadmap for the phased development and enhancement of RAG systems, tailored to meet the evolving demands of industrial applications. Furthermore, we propose knowledge atomizing and knowledge-aware task decomposition to effectively extract multifaceted knowledge from the data chunks and iteratively construct the rationale based on original query and the accumulated knowledge, respectively, showcasing exceptional performance across various benchmarks. The code is publicly available at https://github.com/microsoft/PIKE-RAG.
@@ 5.2 Level-1: Factual Question focused RAG System
5.2.1 Enhanced chunking
@@ 5.3 Level-2: Linkable and Reasoning Question focused RAG System
Chunked text contains multifaceted information, increasing the complexity of retrieval. Recent studies have focused on extracting triple knowledge units from chunked text and constructing knowledge graphs to facilitate efficient information retrieval [edge2024localglobalgraphrag, panda2024holmeshyperrelationalknowledgegraphs]. However, the construction of knowledge graphs is costly, and the inherent knowledge may not always be fully explored. To better present the knowledge embedded the documents, we propose atomizing the original documents in Knowledge Extraction phase, a process we refer as Knowledge Atomizing. Besides, industrial tasks often necessitate multiple pieces of knowledge, implicitly requiring the capability to decompose the original question into several sequential or parallel atomic questions. We refer to this operation as Task Decomposition. By combining the extracted atomic knowledge with the original chunks, we construct an atomic hierarchical knowledge base. Each time we decompose a task, the hierarchical knowledge base provides insights into the available knowledge, enabling knowledge-aware task decomposition.
@@ 5.4 Level-3: Predictive Question focused RAG System
Given the limitations of LLMs in applying specialized reasoning logic, their effectiveness in predictive tasks can be restricted. To overcome this, the knowledge-centric reasoning module is enhanced with a forecasting submodule, enabling the system to infer outcomes based on the input queries and the organized knowledge (e.g. total numbers of medicines approved per year). This forecasting submodule allows the system to not only generate answers based on historical knowledge, but also make projections, providing a more robust and dynamic response to complex queries. By integrating advanced knowledge structuring and prediction capabilities, the L3 system can manage and utilize a more complex and dynamic knowledge base effectively.
@@ 6.3 Evaluation on Legal Benchmarks
Open Australian Legal QA | 25.10 | 34.35 | 63.34 |
@@
Open Australian Legal QA | 16.48 | 88.27 | 98.59 |
@@ 6.4 Real Case Studies
Furthermore, the discrepancy between the formulation of the corpus and the query, is another critical factor advocating for a multi-query approach over a singular deterministic one. The presentation gap can impede the retrieval process even when the generated follow-up question is semantically accurate. For instance, as illustrated in Case (b) in Figure 17, a single-path method such as Self-Ask on the left side might directly inquire ’Who is the mother of Oskar Roehler?’ However, the knowledge base articulates familial relationships using a different schema, ’A is the son of B and C’ in this case, thus the retrieval process falters despite the correctness of the question. Even when we applied the hierarchical retrieval to Self-Ask, the Self-Ask with Hierarchical Retrieval did not succeed in bridging this gap. In contrast, our approach, which generates multiple atomic queries, encompasses a broader range of phrasings that correspond to the diverse representations in the knowledge base. In the depicted case, while the atomic query specifically asking for Oskar Roehler’s mother encounters the same retrieval issue, an alternative query seeking information about his parents successfully retrieves the target chunk. This exemplifies how our method’s flexibility in query generation enhances the likelihood of aligning with the knowledge base’s structure and obtaining accurate information.
@@
Our methodology emphasizes the retrieval of atomic questions rather than directly retrieving chunks. This design choice is exemplified in Case (b) depicted in Figure 17. The knowledge chunk in the corpus is structured using the pattern ’A … as the son of B and C’, which poses challenges for direct retrieval by queries such as ’Who is the mother of …’. In our specialized knowledge base, such direct queries tend to retrieve chunks conforming to the patterns ’A is the mother of B’ or ’A is the father of B’. By utilizing atomic questions as intermediaries for retrieval, our approach effectively narrows the gap between a single query and the multiple sentence structures found in the knowledge base. It facilitates bridging the expression pattern differences exemplified by ’the mother of’ versus ’the son of’ in this scenario.
@@
In contrast to methods like Self-Ask, which only retains intermediate answers for subsequent processing, our method preserves the entire chunk as contextual information. During the atomic selection phase, we present a list of atomic questions as candidate summaries of the relevant content from the original chunk. This strategy significantly reduces token usage and simplifies the process of selecting the pertinent information. Case (c) in Figure 18 demonstrates the dual benefits of our approach: first, by selecting from a curated list of atomic questions, we streamline the identification of relevant information; second, by retaining the entire selected chunk rather than just the intermediate answer, we ensure a rich context is maintained for accurate and comprehensive subsequent processing. While the Self-Ask method on the left retrieves the target chunk, it fails to correctly identify the pertinent ’Ernie Watts’ due to the excessive contextual information. Since retrieved chunks in Self-Ask are discarded after generating an intermediate answer, the method potentially follows an incorrect pathway, leading to an inaccurate conclusion. In contrast, our approach can efficiently filter and select the appropriate atomic question from a concise list. Although the atomic question in this round pertains to the role of Ernie Watts, there is no need to inquire further about his birthplace, as this information is encapsulated within the selected chunk, which remains available for context in subsequent rounds.

### v3

@@ Abstract
Despite notable advancements in Retrieval-Augmented Generation (RAG) systems that expand large language model (LLM) capabilities through external retrieval, these systems often struggle to meet the complex and diverse needs of real-world industrial applications. The reliance on retrieval alone proves insufficient for extracting deep, domain-specific knowledge performing in logical reasoning from specialized corpora. To address this, we introduce sPecIalized KnowledgE and Rationale Augmentation Generation (PIKE-RAG), focusing on extracting, understanding, and applying specialized knowledge, while constructing coherent rationale to incrementally steer LLMs toward accurate responses. Recognizing the diverse challenges of industrial tasks, we introduce a new paradigm that classifies tasks based on their complexity in knowledge extraction and application, allowing for a systematic evaluation of RAG systems’ problem-solving capabilities. This strategic approach offers a roadmap for the phased development and enhancement of RAG systems, tailored to meet the evolving demands of industrial applications. Additionally, we propose knowledge atomizing and knowledge-aware task decomposition to effectively extract multifaceted knowledge from the data chunks and iteratively construct the rationale based on original query and the accumulated knowledge, respectively, showcasing exceptional performance across various benchmarks. Furthermore, we introduce a trainable knowledge-aware decomposer that incorporates domain-specific rationale into the task decomposition and result-seeking process. The code is publicly available at https://github.com/microsoft/PIKE-RAG.
@@ 1 Introduction
We introduce a knowledge-aware task decomposer training strategy that involves collecting rationale-driven data by sampling context and creating diverse interaction. This approach trains the decomposer to incorporate domain-specific rationale into both task decomposition and result-seeking trajectories.
@@ 5.2 Level-1: Factual Question focused RAG System
indicate enhanced chunking and auto-tagging sub-module in knowledge extraction modules.5.2.1 Enhanced chunking
@@ 5.3 Level-2: Linkable and Reasoning Question focused RAG System
indicates atomic knowledge generation in knowledge extraction module, while the squares
represent knowledge reranking and aggregation sub-module in knowledge origination module. Moreover, the squares
indicate the multihop reasoning, comparative reasoning, summarizing sub-modules in knowledge-centric reasoning module.Chunked text contains multifaceted information, increasing the complexity of retrieval. Recent studies have focused on extracting triple knowledge units from chunked text and constructing knowledge graphs to facilitate efficient information retrieval [edge2024localglobalgraphrag, panda2024holmeshyperrelationalknowledgegraphs]. However, the construction of knowledge graphs is costly, and the inherent knowledge may not always be fully explored. To better present the knowledge embedded the documents, we propose atomizing the original documents in Knowledge Extraction phase, a process we refer as Knowledge Atomizing. Besides, industrial tasks often necessitate multiple pieces of knowledge, implicitly requiring the capability to decompose the original question into several sequential or parallel atomic questions. We refer to this operation as Task Decomposition. By combining the extracted atomic knowledge with the original chunks, we construct an atomic hierarchical knowledge base. Each time we decompose a task, the hierarchical knowledge base provides insights into the available knowledge, enabling knowledge-aware task decomposition.
@@ 5.4 Level-3: Predictive Question focused RAG System
indicate knowledge structuring and knowledge induction in knowledge organization module, while the square
represents forecasting sub-module in knowledge-centric reasoning module.Given the limitations of LLMs in applying specialized reasoning logic, their effectiveness in predictive tasks can be restricted. To overcome this, the knowledge-centric reasoning module is enhanced with a forecasting submodule, enabling the system to infer outcomes based on the input queries and the organized knowledge (e.g. total numbers of medicines approved per year). This forecasting submodule allows the system to not only generate answers based on historical knowledge, but also make projections, providing a more robust and dynamic response to complex queries. By integrating advanced knowledge structuring and prediction capabilities, the L3 system can manage and utilize a more complex and dynamic knowledge base effectively.
@@ 6.3 Evaluation on Legal Benchmarks
| Open Australian Legal QA | 25.10 | 34.35 | 63.34 | |
@@
| Open Australian Legal QA | 16.48 | 88.27 | 98.59 | |
@@
6.4 Experiments on Domain Aligned Atomic Proposers
Given that all methods demonstrate notably low accuracy scores on the MuSiQue dataset, we specifically selected this dataset to investigate whether implementing a fine-tuned atomic proposer could potentially enhance overall performance metrics and improve the quality of results.
Data Collection
To establish a comprehensive and efficient experimental framework, we initiate our process by randomly sampling 1000 questions from the dataset. These questions are subsequently divided into two distinct sets following a standard split ratio: 80% of the questions are designated as training data, specifically utilized for collecting and analyzing decomposition trajectories, while the remaining 20% are reserved as an evaluation dataset to assess the model’s performance. For each question within the training dataset, we meticulously follow the procedural framework outlined in Figure 12, systematically generating follow-up questions in a stepwise progression until we successfully obtain a decomposition trajectory that leads to the correct answer. We establish a maximum limit of 10 rounds per question. The results of our data collection process prove to be efficient - we successfully obtain correct answers along with their corresponding decomposition trajectories for 84% of the training data, which represents a substantial improvement compared to the baseline accuracy of 58% achieved without implementing context sampling techniques.
Training Method
Our training methodology is designed to optimize model performance by focusing exclusively on decomposition trajectories that successfully lead to correct answers. The data structure for each record follows a format that captures the complete problem-solving path: , where and are the original question and its answer, respectively. Each pair in the trajectory is composed of an atomic question and its answer. This structured approach ensures that we maintain a clear record of the decomposition process from start to finish.
For our LLM training, we employ supervised fine-tuning (SFT) to enhance model performance. Our process converts each trajectory into structured training data pairs of prompts and responses. These pairs are formatted as according to Algorithm 3. The algorithm incorporates two key functions: and , which generate contextual descriptions based on input parameters. These functions are integral to the transformation process, and we detail their implementation in the following sections.
This transformation is particularly efficient because it generates multiple training instances from each trajectory - specifically, for a trajectory containing sub-questions, we generate distinct training data points, maximizing the learning potential from each successful decomposition.
For all experiments, we use a learning rate of and parameter-efficient fine-tuning (PEFT) with LoRA configuration parameters (lora=16, alpha=64). To ensure reproducibility, we maintain all other hyperparameters at their default values as specified in reference [vonwerra2022trl]. We run all experiments three times with different random seeds and report the mean results. We perform the experiments on a computing node equipped with a single NVIDIA A100-80G GPU for all training processes.
Experiment Results
Table 11 presents our experimental findings. We select three open-source LLMs with varying parameter sizes and computational capabilities as our base models for atomic proposers: meta-llama/Llama-3.1-8B [llama3modelcard], Qwen/Qwen2.5-14B [qwen2], and microsoft/phi-4 [abdin2024phi4technicalreport]. Through SFT, we develop these into fine-tuned atomic proposers. For inference, we integrated two advanced LLMs-GPT-4o (version 2024-11-20) and Llama-3.1-70B-Instruct [llama3modelcard]-to process the retrieved data and generate the final answers. These models were chosen for their enhanced capabilities in delivering accurate, high-quality outputs.
For illustration, the second row and second column in Table 11 shows our use of the original meta-llama/Llama-3.1-8B as the atomic proposer for question decomposition. After gathering sufficient context, we use GPT-4o to generate the final answer. In contrast, the second row and third column shows results when using a fine-tuned version of meta-llama/Llama-3.1-8B as the atomic proposer, while still using GPT-4o for final answer generation. The difference in values demonstrates how fine-tuning atomic proposers with domain-specific decomposition trajectories improves performance.
| GPT-4o | GPT-4o+FT | Llama-3.1-70B-Instruct | Llama-3.1-70B-Instruct+FT | |
| Llama-3.1-8B | 47.83% | 62.14% | 48.37% | 58.70% |
| Qwen2.5-14B | 56.52% | 63.95% | 57.61% | 63.04% |
| phi-4-14B | 60.33% | 65.76% | 58.70% | 62.50% |
@@ 6.5 Real Case Studies
Furthermore, the discrepancy between the formulation of the corpus and the query, is another critical factor advocating for a multi-query approach over a singular deterministic one. The presentation gap can impede the retrieval process even when the generated follow-up question is semantically accurate. For instance, as illustrated in Case (b) in Figure 17, a single-path method such as Self-Ask on the left side might directly inquire ‘Who is the mother of Oskar Roehler?’ However, the knowledge base articulates familial relationships using a different schema, ‘A is the son of B and C’ in this case, thus the retrieval process falters despite the correctness of the question. Even when we applied the hierarchical retrieval to Self-Ask, the Self-Ask with Hierarchical Retrieval did not succeed in bridging this gap. In contrast, our approach, which generates multiple atomic queries, encompasses a broader range of phrasings that correspond to the diverse representations in the knowledge base. In the depicted case, while the atomic query specifically asking for Oskar Roehler’s mother encounters the same retrieval issue, an alternative query seeking information about his parents successfully retrieves the target chunk. This exemplifies how our method’s flexibility in query generation enhances the likelihood of aligning with the knowledge base’s structure and obtaining accurate information.
@@
Our methodology emphasizes the retrieval of atomic questions rather than directly retrieving chunks. This design choice is exemplified in Case (b) depicted in Figure 17. The knowledge chunk in the corpus is structured using the pattern ‘A … as the son of B and C’, which poses challenges for direct retrieval by queries such as ‘Who is the mother of …’. In our specialized knowledge base, such direct queries tend to retrieve chunks conforming to the patterns ‘A is the mother of B’ or ‘A is the father of B’. By utilizing atomic questions as intermediaries for retrieval, our approach effectively narrows the gap between a single query and the multiple sentence structures found in the knowledge base. It facilitates bridging the expression pattern differences exemplified by ‘the mother of’ versus ‘the son of’ in this scenario.
@@
In contrast to methods like Self-Ask, which only retains intermediate answers for subsequent processing, our method preserves the entire chunk as contextual information. During the atomic selection phase, we present a list of atomic questions as candidate summaries of the relevant content from the original chunk. This strategy significantly reduces token usage and simplifies the process of selecting the pertinent information. Case (c) in Figure 18 demonstrates the dual benefits of our approach: first, by selecting from a curated list of atomic questions, we streamline the identification of relevant information; second, by retaining the entire selected chunk rather than just the intermediate answer, we ensure a rich context is maintained for accurate and comprehensive subsequent processing. While the Self-Ask method on the left retrieves the target chunk, it fails to correctly identify the pertinent ‘Ernie Watts’ due to the excessive contextual information. Since retrieved chunks in Self-Ask are discarded after generating an intermediate answer, the method potentially follows an incorrect pathway, leading to an inaccurate conclusion. In contrast, our approach can efficiently filter and select the appropriate atomic question from a concise list. Although the atomic question in this round pertains to the role of Ernie Watts, there is no need to inquire further about his birthplace, as this information is encapsulated within the selected chunk, which remains available for context in subsequent rounds.